# Common Corruptions for Enhancing and Evaluating Robustness in Air-to-Air Visual Object Detection

### v1

@@ I Introduction
Yet, the scenarios encompassed by these publicly available benchmarks are typically constrained. Notably, these datasets often lack representations of adverse weather effects, despite the potential impact of fog or rain on images as elucidated in studies [10, 11]. Beyond external scenarios, the internal noise inherent in sensors can further contribute to increased deviation and variance in ranging measurements. Ranging measurements, as highlighted in studies [12, 13], can lead to data corruption and a subsequent decline in detector performance due to internal noise in the sensors. Therefore, it is imperative to conduct a thorough evaluation of an object detector under these varied corruptions before deploying it in real-world environments
@@
  • We demonstrate that fine-tuning our synthetically constructed corruptions improves the performance in real-world flight experiments that include challenging conditions and significantly differ from the training dataset (domain shift)
@@ II-B Robustness Benchmarks
Several studies [14, 19, 20] [21, 22, 23] have demonstrated that Deep Neural Networks (DNNs) exhibit susceptibility to common corruptions. To methodically explore and enhance DNN robustness against corruptions, corruption benchmarks were initially introduced in image recognition [24] [25, 26, 27] and later extended to 3D object detection [12], semantic segmentation [28], pose estimation [29], and person Re-ID [30].
However, numerous investigated corruptions are hypothetical and may not accurately represent real-world scenarios [31, 32, 33, 34, 35, 36, 37] in autonomous navigation of UAVs. Moreover, in the context of air-to-air aerial detection, the objects are typically of small size. Hence, the artificially generated corruptions need to be crafted in a manner that ensures the object remains visible across all severity levels. As a consequence, developing a comprehensive benchmark for evaluating the robustness of air-to-air aerial object detection across diverse real-world flying conditions remains a challenging task [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. To the best of our knowledge, in this work, we introduce the first robustness benchmark dataset for aerial object detection.
@@ III-B Discussion on the disparity between synthetic and real-world corruptions
Additionally, each corruption type exhibits four severity levels, representing different intensities of manifestation. An example depicting four severity levels for each type of our synthesized corruptions is illustrated in Fig. 1. These corruptions are implemented using functions, enabling seamless integration into the data loader for enhanced portability and storage efficiency
@@ IV-A Benchmark Dataset
All detectors are executed based on the open-source codes released on GitHub. For YOLOv5, YOLOv8 and YOLOX detectors we chose the large type model because is more suitable for small object detection [61]. The training and evaluation are all executed on the NVIDIA RTX 4090 GPU with a memory of 24GB. The batch size of each detector is optimized to reach the limit of GPU memory.

### v2

# Common Corruptions for Evaluating and Enhancing Robustness in Air-to-Air Visual Object Detection

@@ I Introduction
Yet, the scenarios encompassed by these publicly available benchmarks are typically constrained. Notably, these datasets often lack representations of adverse weather effects, despite the potential impact of fog or rain on images as elucidated in studies [10, 11]. Beyond external scenarios, the internal noise inherent in sensors can further contribute to increased deviation and variance in ranging measurements. Ranging measurements, as highlighted in studies [12, 13], can lead to data corruption and a subsequent decline in detector performance due to internal noise in the sensors. Therefore, it is imperative to conduct a thorough evaluation of an object detector under these varied corruptions before deploying it in real-world environments.
@@
  • We demonstrate that fine-tuning our synthetically constructed corruptions improves the performance in real-world flight experiments that include challenging conditions and significantly differ from the training dataset (domain shift).
@@ II-B Robustness Benchmarks
Several studies [14, 19, 20, 21] have demonstrated that Deep Neural Networks (DNNs) exhibit susceptibility to common corruptions. Especially, [21] suggested that detecting objects irrespective of image distortions or weather conditions is essential for practical implementations of deep learning, such as autonomous driving. To methodically explore and enhance DNN robustness against corruptions, corruption benchmarks were initially introduced in image recognition [22] and later extended to 3D object detection [12], semantic segmentation [23], pose estimation [24], and person Re-ID [25]. In addition, simulated imagery is used for air-to-ground object detection in [26, 27].
However, numerous investigated corruptions are hypothetical and may not accurately represent real-world scenarios in autonomous navigation of UAVs. Moreover, in the context of air-to-air aerial detection, the objects are typically of small size. Hence, the artificially generated corruptions need to be crafted in a manner that ensures the object remains visible across all severity levels. As a consequence, developing a comprehensive benchmark for evaluating the robustness of air-to-air aerial object detection across diverse real-world flying conditions remains a challenging task. To the best of our knowledge, in this work, we introduce the first robustness benchmark dataset for aerial object detection.
@@ III-B Discussion on the disparity between synthetic and real-world corruptions
Additionally, each corruption type exhibits four severity levels, representing different intensities of manifestation. An example depicting four severity levels for each type of our synthesized corruptions is illustrated in Fig. 1. These corruptions are implemented using functions, enabling seamless integration into the data loader for enhanced portability and storage efficiency.
@@ IV-A Benchmark Dataset
All detectors are executed based on the open-source codes released on GitHub. For YOLOv5, YOLOv8 and YOLOX detectors we chose the large type model because is more suitable for small object detection [37]. The training and evaluation are all executed on the NVIDIA RTX 4090 GPU with a memory of 24GB. The batch size of each detector is optimized to reach the limit of GPU memory. The default optimizers are employed, and parameters such as learning rate (LR), momentum, weight decay and batch size are meticulously adjusted through extensive experimentation.
@@ IV-B Real-world Flight tests
To provide further context, the real-flight dataset employed in our study encompasses $10,641$ images captured from two distinct drones equipped with different cameras. The first drone, a DJI Mavic, utilized its onboard camera to capture images at a resolution of $4K(3840\times 2160)$ pixels, operating at a frame rate of $30$ frames per second (fps). The second drone, a custom model with approximate dimensions of $1000\times 1000\times 500$ mm ($width\times depth\times height$), utilized a GoPro Hero $8$ camera, capturing images at a resolution of $4000\times 3000$ pixels, with a frame rate of $25$ fps. Both camera streams were processed at a uniform frame rate of $5$ fps to ensure consistency in data processing.